Assume the following information for Halloween Corp.

Accounts receivable (beginning balance) $143,000
Allowance for doubtful accounts (beginning balance) 11,470
Net credit sales 950,000
Collections 902,000
Write-offs of accounts receivable 5,500
Collections of accounts previously written off 2,300

Uncollectible accounts are expected to be 9% of the ending balance in accounts receivable.

a) Prepare the entries to record sales and collections during the period, a separate entry to record the write-off of uncollectible accounts during the period, and another separate entry to record bad debt expense for the period.

b) Prepare the entries to record the recovery of the uncollectible account during the period.

c) Determine the ending balance in Accounts Receivable and the unadjusted balance in Allowance for Doubtful Accounts.

d) Determine the ending (adjusted) balance in Allowance for Doubtful Accounts.
Answer #1

a) Journal entries

Date account and explanation Debit Credit
Account receivable 950000
Sales revenue 950000
(To record sales)
Cash 902000
Account receivable 902000
(To record collection)
Allowance for doubtful accounts 5500
Account receivable 5500
(To record write off)
Bad debt expense (185500*9%-8270) 8425
Allowance for doubtful accounts 8425
(To record bad debt expense)

Journal entry

date account and explanation Debit Credit
Account receivable 2300
Allowance for doubtful accounts 2300
(To record reinstate)
Cash 2300
Account receivable 2300
(To record collection)

c) Ending balance of account receivable = 143000+950000-902000-5500 = 185500

Unadjusted balance for doubtful accounts = 11470-5500+2300 = 8270

d) Ending balance in allowance for doubtful accounts = 185500*9% = 16695